OpenStudy (anonymous):

If xy/x+y=1 what is x

OpenStudy (anonymous):

\[\frac{xy}{x+y}=1\iff xy=x+y\]

OpenStudy (anonymous):

i need to find x not xy

OpenStudy (anonymous):

Yes, satellite is doing it in the right way. Find two numbers whose product is also equal to their sum.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

your answer will have a \(y\) in it i believe a one answer could be both are 0 which you can see by inspection another you might try is both are 2

OpenStudy (anonymous):

if you start with \[xy=x+y\] you can solve for \(x\) by writing \[xy-x=y\] \[x(y-1)=y\] \[x=\frac{y}{y-1}\] so if you pick any \(y\) you can find the corresponding \(x\)

OpenStudy (anonymous):

well these are the choices: a) y+1/y-1 b) y+1/y c) y/y-1 d) y/y+1 e) 1-1/y

OpenStudy (anonymous):

ok then use the answer above

OpenStudy (anonymous):

or just go with C because it is almost always C in fact C is right in this case (as usual)